HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

EXPLORING THE DIFFERENCES IN ORTHO- AND PARA-H_{2}-FLUOROBENZALDEHYDE COMPLEXES USING FTMW METHODS

\underline{B. KEMPKEN} , R. M. DOHMEN, D. A. OBENCHAIN, Institute of Physical Chemistry, Tammannstrasse 6, Georg-August-Universität Göttingen, 37077 Göttingen; P. PINACHO, M. SCHNELL, Deutsches Elektronen-Synchrotron

Notkestrasse 85, 22607 Hamburg

Understanding the intermolecular binding of molecular hydrogen to other compounds is a basic requirement for the design of possible H_{2} storage materials. Nevertheless, studies on H_{2} van der Waals complexes with e.g. H_{2}O,^{1} OCS^{2}, HCN^{3}, and AgCl^{4} show that to date, there are still gaps in the comprehension especially concerning differences in the binding of ortho- and para-H_{2} to a specific host molecule. Our group investigates the structure of bimolecular ortho- and para-H_{2} complexes with 2-, 3-, and 4-fluorobenzaldehyde, applying chirped pulse and high-resolution cavity Fourier-transform microwave (FTMW) spectroscopy. Rotational and centrifugal distortion constants were determined, supported by DFT-based predictions. Furthermore, the cavity FTMW measurements, taken with a Fabry-Pérot spectrometer, enabled us to resolve the nuclear spin-spin interaction for complexed ortho-H_{2}. The measured rotational constants were used to fit the intermolecular distance. We found that the H_{2}’s center of mass distance to the aromatic plane is 0.36 – 0.37 nm for ortho-H_{2} and 0.34 – 0.35 nm for para-H_{2} complexes. This result is surprising since the mentioned studies encourage the conclusion that ortho-H_{2} generally binds more strongly to a host molecule than para-H_{2}. A possible explanation could be a coupling distortion of the ortho-H_{2}’s rotational energy levels due to internal rotation of the H_{2} group, which we investigate here. DFT-predicted rotational barriers are between 16 and 32 cm^{−1}.

1doi:10.1063/S0009-2614(03)00476-7, X.-G. Wang, T. Carrington, Theoretical study of the rovibrational spectrum of H_{2}O-H_{2}, J. Chem. Phys. 134, 044313 (2011).

2doi:10.1080/00268970210138904, Z. Yu, K. J. Higgins, W. Klemperer, M. C. McCarthy, P. Thaddeus, K. Liao, W. Jäger, Rotational spectra of the van der Waals complexes of molecular hydrogen and OCS, J. Chem. Phys. 127, 054305 (2007).

3doi:10.1063/1.462117, M. Ishiguro, T. Tanaka, K. Harada, C. J. Whitham, K. Tanaka, Rotational spectra of the H_{2}-HCN cluster observed by millimeter-wave spectroscopy combined with a pulsed supersonic jet technique, J. Chem. Phys. 115, 5155 (2001).

4doi:10.1364/AO.36.008533, G. S. Grubbs II, D. A. Obenchain, H. M. Pickett, S. E. Novick, H_{2}- AgCl: A spectroscopic study of a dihydrogen complex, J. Chem. Phys. 141, 114306 (2014).

p-number: p206  Submitted on Fri, 16 Jun 2023 10:13:49 +0200